OpenStudy (anonymous):

witch would be an example of associative: 1) (3 + 6x) - 5x 2)3 + (6x - 5x) 3)3 + (6 - 5)x 4)3 + (1)x 5)3 + x

OpenStudy (bloomlocke367):

When you have the associative property, you just change the grouping, but you would still get the same answer.

OpenStudy (anonymous):

oh ok but what about multiplicative identity

OpenStudy (anonymous):

commutative

OpenStudy (anonymous):

and distributive

OpenStudy (anonymous):

i'm not good with math

OpenStudy (bloomlocke367):

okay, multiplicative identity means you can multiply any number by 1, and you'll get the same number. and it's okay you're not good with math. Some people are, others aren't c: it's not a problem.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:)

OpenStudy (bloomlocke367):

the commutative property is when you can change the order of numbers and you'll end up with the same thing. This would only work with addition or multiplication. For example 'a+b=b+a'. Does this stuff make sense to you?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

it helps me for a test i'm studying for

OpenStudy (anonymous):

at miner acadamy at bauxite high school

OpenStudy (bloomlocke367):

You're welcome. :) The distributive property allows you to multiply a number or term with more than one term inside a set of parentheses. Would you like an example of that?

OpenStudy (anonymous):

yes please

OpenStudy (bloomlocke367):

 You would distribute the 3 to both the 2x and the 6 and multiply them.

OpenStudy (bloomlocke367):

Sorry that looks bad XD

OpenStudy (texaschic101):

my brain must be fried because I am not seeing associative property in your answer choices....I see distributive, but not associative.

OpenStudy (anonymous):

it's ok

OpenStudy (texaschic101):

(3 + 6x ) - 5x = 3 + (6x - 5x)......that would be associative

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

possibly 3 as your answer... 3 + (6-5)x = 3 + 6x - 5x.....but I think that is distributive

OpenStudy (texaschic101):

read this...it should help http://www.purplemath.com/modules/numbprop.htm

OpenStudy (bloomlocke367):

The one you gave is distributive, @texaschic101

OpenStudy (bloomlocke367):

because you'd have to distribute the x

OpenStudy (bloomlocke367):

I, personally, think it's option 2, because 6x-5x can be grouped without the final solution being changed.

OpenStudy (texaschic101):

but look at the link I posted..look under associative...it has the x following the parenthesis as an example of associative

OpenStudy (bloomlocke367):

Well, the way you solved it... made it distributive. it's all about how you solve it in some cases XD